OpenStudy (anonymous):

A boy and a girl have masses of 45 kg and 32 kg respectively. Both are balanced on the opposite ends of a 5.0 m long wooden plank with a mass of 16 kg. At what point along the plank does the pivot point have to be?

OpenStudy (anonymous):

 here the plank is in equilibrium only if the torques on both sides are equal .but the problem is with the mass of the plank . let the plank be having a mass per unit length "p" then p=16/5. now we can find torques on each side of the pivot on the left side torque = 45*x + x/2 * (x*p) {****** this is the torque by the planks weight , the plank to the left has a wight of p*x and this weight acts at midway x/2} similarly on the right side torque =32*(5-x) + (5-x)/2 * (p* 5-x) now solve for x